# EgoHand: Ego-centric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMUs

### v1

@@ 5 Conclusion
In this paper, we present EgoHand, an proof-of-concept solution for VR interaction gesture recognition utilizing millimeter-wave radar and IMUs worn on the head for preserving more personal privacy. We conduct extensive cross-domain and few-shot experiments to evaluate EgoHand, highlighting the importance of training with samples from both hands to enhance cross-hand stability. Additionally, at least two pre-use calibrations are essential to ensure optimal performance across various users, scenes, and postures.
Current EgoHand appears somewhat cumbersome, particularly with the TI mmWave radar. This can potentially be improved, for example, by using Google’s Soli radar (9mm×$\times$9mm) [23]. In addition, mmWave radar may expose health metrics like breathing and heart rate [44, 4], which can be mitigated by adjusting the range and angle of radar signals.

### v2

@@ 5 Conclusion
This paper presents mmEgoHand, a proof-of-concept system for egocentric hand pose estimation and gesture recognition, leveraging head-mounted millimeter-wave radar and IMUs. This configuration facilitates user mobility while offering enhanced personal privacy compared to traditional camera-based solutions. We conduct extensive experiments to evaluate mmEgoHand, demonstrating the importance of including samples from both hands during training to improve cross-hand generalization. Furthermore, we find that performing at least two pre-use calibration sessions is critical to achieving robust performance across diverse users, environments, and hand postures.
While mmEgoHand  shows promising results, the current prototype is somewhat cumbersome, especially due to the use of the TI mmWave radar module. Future versions could benefit from more compact hardware, such as Google’s Soli radar (9mm×$\times$9mm) [50]. Additionally, mmWave radar has the potential to reveal health-related metrics such as respiration and heart rate [51, 52], which raises privacy concerns. These can be mitigated by appropriately limiting the radar’s range and field of view.
Moreover, although mmEgoHand  achieves promising accuracy in gesture recognition, there remains room for improvement in the precision of hand pose estimation and its generalization under cross-domain scenarios. To facilitate further research and development, we have publicly released our dataset, enabling the community to explore, benchmark, and improve upon our work.